# TAVP: Task-Adaptive Visual Prompt for Cross-domain Few-shot Segmentation

### v1

@@ V-B Ablation Study
Branch Ablation. In branch ablation testing based on different backbones, we chose ResNet as the backbone of CNN-based feature extraction in CDTAP. First, MFF is multi-level feature fusion module. CDTAP is a task-adaptive information disentanglement module. Previous work of APSeg [50] and HQ-SAM [36] was set along with our method for branch ablation study, shown in Table III.
TABLE III: Ablation Study of MFF and CDTAP in 5-shot setting.
@@
APSeg ViT 86.91 69.14 42.03 82.50 70.14
Our ViT+ResNet 88.61 73.39 61.98 83.41 76.85
Comparition with SOTA under the Same Setting. To provide a fair baseline, we train PATNet with ViT-base, SAM initialization, and $1024\times 1024$ crops. Results in Table IV demonstrate again the superiority of our TAVP compared with PATNet under the same settings. We can see that the ViT-based PATNet improves the performance compared to CNN-based PATNet [4] on the dataset of Chest X-ray, ISIC, and Deepglobe. These results prove again that these three datasets are more challenging, and our method performs more robustly.
@@
V-C Efficiency Comparision
Considering the huge amount of parameter calculation required for the basic model, we only train some fine-tunable parameters. SAM needs to train a model with a large number of parameters from scratch, while our framework only needs to fine-tune some layers and parameters instead of starting from scratch. In addition, linear computation is added to our framework to reduce the number of parameters, which results in our model being more parametric lightweight compared with SAM.
@@ V-C Efficiency Comparision
The detailed efficiency comparison results are shown in the Table VI. Table V shows the detailed parameter comparison. Notice that the bottom three lines are the parameters of the original SAM based on ViT. The other lines are our whole framework’s parameters based on different backbones, all smaller than SAM. These results prove again that our method improves the efficiency of SAM and is more light-weighted.
VI Conclusion
It is worth noting that our work is the first one to apply a large foundation model-based method to CDFSS tasks. Previous works have focused on traditional CNN-based deep learning methods, slowing down the exploration of Artificial General Intelligence. Our method is based on SAM, a large foundational model for segmentation, concurrently rethinking its value in CDFSS. Extensive experiments have shown that although SAM can get satisfactory results for most segmentation tasks, it still performs worse in some particular scenarios, especially on the Deepglobe data set. Thus, our work has provided a novel and efficient framework for large model transfer in CDFSS tasks. The CDTAP module achieves learnable prompts for excellent performance on the three benchmarks.
In other words, SAM can be a basic knowledge tool, and using SAM to transfer its knowledge to other specific situations to complete domain-specific tasks is worthwhile. Besides, our method is an initial exploration of transferring SAM to CDFSS tasks, and many of the more critical and efficient algorithms need to be explored in the future. More methods with strong learning ability that can complete cross-domain and few-shot tasks need to be researched, helping the development of Artificial General Intelligence in the future.

### v2

@@ IV-C Ablation Study
Backbone and Data Augmentation Ablation. In the additional task-specific class-domain agnostic feature extraction and auto-prompts generation branch, we perform ablation experiments on the models’ performance on three datasets that are more difficult for cross-domain challenges under 1-way 1-shot and 5-shot settings, and the details are shown in Table II. This study proves that ResNet has a stronger recognition ability for categories, and its effect is outstanding in our novel learnable prompt.
Branch Ablation. In branch ablation testing based on different backbones, we chose ResNet as the backbone of CNN-based feature extraction in CDTAP. First, MFF is a multi-level feature fusion module. CDTAP is a task-adaptive information disentanglement module. We performed ablation experiments with our method in the same setting as previous work of APSeg [61] and HQ-SAM [44], and the results are description in Table III.
TABLE III: Ablation Study of MFF and CDTAP in the 5-shot setting.
@@
APSeg ViT 86.91 71.14 47.63 83.41 72.27
Our ViT + ResNet 87.03 73.39 60.98 82.90 76.08
Comparison with SOTA under the Same Setting. To provide a fair baseline, we train PATNet with ViT-base, SAM initialization, and $1024\times 1024$ crops. Results in Table IV demonstrate again the superiority of our TAVP compared with PATNet under the same settings. We can see that the ViT-based PATNet improves the performance compared to CNN-based PATNet [6] on the dataset of Chest X-ray, ISIC, and Deepglobe. These results prove again that these three datasets are more challenging, and our method performs more robustly.
@@
IV-D Efficiency Comparison
Considering the huge amount of parameter calculation required for the basic model, we only train some fine-tunable parameters. SAM needs to train a model with a large number of parameters from scratch, while our framework only needs to fine-tune some layers and parameters instead of starting from scratch. In addition, linear computation is incorporated into our framework to reduce the number of parameters, thus requiring significantly fewer parameters.
@@ IV-D Efficiency Comparison
The detailed comparison results of efficiency are shown in the Table VI. Table V shows the detailed parameter comparison. Notice that the bottom three lines are the parameters of the original SAM based on ViT. The other lines are our whole framework’s parameters based on different backbones, all smaller than SAM. These results prove again that our method improves the efficiency of SAM and is more light.
IV-E T-SNE Visualization
We use t-SNE plots to visualize the distribution of test data from a specific target domain. In Figure 5 (a), the original distribution shows that the foreground-boundary (F-B) and background-boundary (B-B) pixels are mixed together, making it difficult to clearly differentiate between the foreground and background regions. In contrast, (b) demonstrates the effect of applying a prototype-based clustering method, which reorganizes the data and allows for a more distinct separation of foreground and background information. This shows how t-SNE visualization, combined with prototype-based clustering, enhances the clarity and structure of the data distribution.
Figure 6: Visualization of foreground-boundary (F-B) and background-boundary (B-B) pixels in t-SNE plots. (a) The original distribution showing the F-B and B-B pixels. (b) The prototype-based distribution displaying the F-B and B-B pixels based on prototype-based clustering.
Figure 7: Performance of our model on a medical image dataset: Lung-CT, the 2d-predictions is shown in (c). We test the 3-dimensional data and print the test mask using a visualization tool, as shown in (e).
IV-F Performance on other datasets
Except for the benchmarks we compared on the paper, we also tested on other few-shot and cross-domain datasets including CT-Lung, and SUIM.
Performance on CT-Lung [62] The dataset derived from the Lung Nodule Analysis (LUNA) competition is a collection of CT scans focused on the lungs. The previous study [63] used this dataset to test cross-domain performance. It encompasses 534 CT images of the lung, each with a resolution of 512 by 512 pixels. Notably, this dataset is exclusively dedicated to lung-related imagery, and all the images within it are in grayscale. We tested the model combining CDFS branch and auto-prompt branch, the visualization of performance is shown in Figure  7.
Figure 8: Performance of our model on an underwater image dataset: SUIM, the predictions are shown in (c) and the overlay is visualized in blue color.
Performance on SUIM [64] The SUIM dataset is a specialized resource designed for underwater image segmentation. It comprises a collection of 1,525 images, featuring eight distinct classes: fish, reefs, aquatic plants, wrecks/ruins, human divers, robots, and sea-floor. The dataset is tailored to enhance the performance and accuracy of image segmentation algorithms in underwater environments, a challenging domain due to factors like varying light conditions and water turbidity. By offering a variety of underwater scenes and objects, SUIM plays a crucial role in advancing computer vision technologies for marine research, underwater robotics, and environmental monitoring. We tested the model combining CDFS branch and auto-prompt branch, the visualization of performance is shown in Figure  8.
V Conclusion
It is worth noting that our work is the first one to apply a large foundation model-based method to CD-FSS tasks, shifting focus from traditional CNN-based deep learning approaches. By leveraging the Segment Anything Model (SAM), a powerful foundational model for segmentation, the proposed framework redefines SAM’s role in CD-FSS tasks and introduces a novel perspective on using large models to address domain-specific challenges. The incorporation of the CDTAP module, which enables adaptive and learnable visual prompts, allows for enhanced segmentation accuracy and robustness, achieving state-of-the-art performance on three widely-used CD-FSS benchmarks.
The extensive experiments conducted demonstrate that SAM provides satisfactory results for a variety of segmentation tasks, showcasing its generalization capability. However, the study also highlights limitations in certain scenarios, such as the DeepGlobe dataset, where SAM’s performance does not meet expectations, underlining the necessity for further refinement of SAM-based methods to enhance their adaptability and effectiveness in more challenging environments. The proposed framework thus serves as a significant step forward, offering an innovative and efficient pathway for large model transfer in CD-FSS tasks. Beyond practical outcomes, this work opens a new frontier in leveraging foundational models for cross-domain and few-shot learning. SAM’s ability to act as a foundational knowledge tool, transferring its learned representations to new and diverse tasks, is a noteworthy achievement. The adaptive visual prompts introduced in this study provide a flexible mechanism for knowledge transfer, demonstrating the potential for SAM-based approaches to tackle domain-specific segmentation problems effectively.
Moreover, this work serves as an initial exploration into the transfer of SAM’s knowledge through adaptive visual prompts, emphasizing the need for future research into more efficient algorithms with strong learning capabilities. Such advancements will not only enhance CD-FSS performance but also contribute to the broader goal of advancing Artificial General Intelligence by enabling more robust domain adaptation and few-shot learning methodologies.